U.S. flag

An official website of the United States government

Dot gov

Official websites use .gov
A .gov website belongs to an official government organization in the United States.

HTTPS

Secure .gov websites use HTTPS
A lock () or https:// means you’ve safely connected to the .gov website. Share sensitive information only on official, secure websites.

Layline Petroleum Oil Spill-FPN E12608

All POL/SITREP's for this site Layline Petroleum Oil Spill-FPN E12608
Breckenridge, TX - EPA Region VI
POLREP #5
Final
Printer Friendly  |   PDF
 
U.S. ENVIRONMENTAL PROTECTION AGENCY
POLLUTION/SITUATION REPORT
Layline Petroleum Oil Spill-FPN E12608 - Removal Polrep
Final Removal Polrep

EPA Emergency Response

UNITED STATES ENVIRONMENTAL PROTECTION AGENCY
Region VI

Subject: POLREP #5
Final
Layline Petroleum Oil Spill-FPN E12608
FPN E12608
Breckenridge, TX
Latitude: 32.9007600 Longitude: -98.8010200


To:
From: Nicolas Brescia, OSC
Date: 12/21/2011
Reporting Period: 12/17/2011 to 12/20/2011

1. Introduction
  1.1 Background
   
Site Number: FPN E12608    Contract Number:  
D.O. Number:      Action Memo Date:  
Response Authority: OPA    Response Type: Emergency
Response Lead: PRP    Incident Category:
NPL Status:    Operable Unit:
Mobilization Date:      Start Date:  
Demob Date:      Completion Date:  
CERCLIS ID:    RCRIS ID:
ERNS No.:    State Notification:
FPN#: E12608    Reimbursable Account #:

1.1.1 Incident Category
Crude oil/brine spill.

1.1.2 Site Description
The release occurred at the A.A. Atkins Lease (03900).

1.1.2.1 Location
The lease is located in Stephens County off of County Road 176.

1.1.2.2 Description of Threat
A 3 inch poly high pressure gathering line breached releasing an estimated 200 BBL of crude oil and 1000 BBL of brine into a ravine. The oil/salt water migrated from the ravine into Long Branch Creek. Long Branch Creek flows into Bufford Creek which flows into the Clear Fork Brazos River.



1.1.3 Preliminary Removal Assessment/Removal Site Inspection Results
On 12/09/2011 EPA OSC Brescia and EPA START arrived on scene, received site access, and conducted assessment activities. OSC Brescia met with the RP's foreman and conducted a site walk from the release area to Long Branch Creek. EPA START arrived on scene and conducted a complete site walk and assessment of the impacted creek area.

2. Current Activities
  2.1 Operations Section
    2.1.1 Narrative
On 12/06/2011 Layline Petroleum (LP) discovered a break in a 3 inch poly high pressure gathering line which released approximately 200 BBL of crude oil and 1000 BBL of salt water into Long Branch Creek. LP contacted the Texas Railroad Commission (TXRRC) at 1400. The TXRRC sent a field investigator to the site. TXRRC notified the Texas Parks and Wildlife(TXPW). TXPW mobilized a spill biologist to the scene to conduct an assessment.

On 12/09/2011, TXPW notified EPA of the spill. EPA OSC Rhotenberry contacted the Responsible Party (LP) and informed them that they were required to contact the National Response Center(NRC) to report the release since a waterway was impacted. LP then made the report to the NRC, NRC Report 997681. EPA OSC Rhotenberry dispatched EPA OSC Brescia and EPA START to the site to conduct an assessment.

2.1.2 Response Actions to Date
On 12/09/2011 at approximately 1400, EPA OSC and EPA START arrived on site and conducted site assessment activities which included documenting impacts to the creek and determining flow from Long Branch Creek. During the assessment activities, OSC Brescia met with the RP's foreman and discussed EPA's activities. Layline Petroleum (LP) foreman explained to OSC Brescia that he had contractors on site, that they had currently stopped flow of the oil in the creek, and that they were currently recovering pockets of oil and salt water utilizing vacuum trucks.

EPA assessment activities found that Long Branch Creek is classified as an intermittent stream. The creek currently held water. Long Branch Creek flows into Bufford Creek which flows into the Clear Fork Brazos River. The entrance of the oil into the creek is located approximately 1.3 miles from Bufford Creek, and approximately 2.3 miles from the Clear Fork Brazos River. A potential threat existed for the oil to migrate from the creek into Bufford Creek and then into the Clear Fork Brazos River. OSC Brescia discussed further response actions by LP and LP's foreman informed OSC Brescia that they would continue recovery and cleanup efforts as required. LP had approximately 12 personnel on site utilizing three vacuum trucks and one land based pump to recover oil from the creek. LP was also retaining additional vacuum trucks and personnel.

Upon completion of the initial site assessment conducted by EPA, it was determined that approximately 0.5 miles of the creek had been impacted, and that approximately six large pools of oil existed within the creek. Sandstone rock which lined the creek area had been oiled along with vegetation and pockets of oil was present on sediment. Assessment activities did not document any wildlife or fish affected from the release. TXPW did discover several dead fish impacted from the release. OSC Brescia contacted TXPW and requested any information on any endangered/threatened species in the creek. TXPW informed OSC Brescia that none were present. LP's foreman informed OSC Brescia that LP would continue operations until the creek was clean. OSC Brescia discussed a verbal scope of work with the LP's foreman and would provide a written scope of work to LP on 12/10/2011. OSC Brescia requested that LP have underflow dam materials on site in case a rain event occurred to minimize any oil released from the creek. LP informed OSC Brescia that they would obtain the materials.

On 12/10/2011, OSC Brescia met with the RP's foreman and presented a written scope of work for cleanup activities. LP's foreman agreed to the scope of work and signed copies were provided to all parties. The scope of work included the following: recover all oil from the creek area utilizing mechanical and manual techniques, removal of pockets of oil on sediment utilizing mechanical or manual techniques, flush impacted areas with water and recover rinsate water, removal of all heavily oiled vegetation from the creek, and removal of oil impacted soil from the release point to the creek entrance. EPA START continued documentation of the cleanup efforts. LP continued oil recovery operations utilizing three vacuum trucks and began sediment removal utilizing shovels. LP also began water flushing activities utilizing water trucks at the point of entry into the creek working down the creek bank. LP created a berm area utilizing soil and poly to place the impacted soil from the ravine. Since recovery operations began on 12/07/2011, approximately 2,000 BBL of oil/salt water/water have been collected from two pockets in the creek. Out of the 2,000 BBL recovered, an estimated 50 BBL is oil. An additional crew of 8 personnel arrived on site to assist in the recovery and removal efforts. Once it was determined that LP was conducting appropriate cleanup activities, OSC Brescia and EPA START demobilized from the scene. EPA START will return on 12/12/2011 to continue documenting cleanup activities. 

During the period of 12/11/2011 through 12/12/2011, LP continued recovery operations focusing on the remaining four pools of oil. All pools of oil were recovered utilizing vacuum trucks. During the recovery operations, an additional 1,600 BBL of oil/salt water/water was recovered. Out of the 1,600 BBL recovered, an estimated 20 BBL is oil. Oiled sediment and vegetation was also collected along the creek. LP continued flushing operations and have currently flushed approximately 0.17 miles of the 0.5 miles of impacted creek area. LP excavated approximately 90% of the oil impacted soil from the source to the creek entrance, and they staged the soil in a bermed poly area. LP continues to work 24/hrs a day with response contractors to complete the SOW. EPA START was on site documenting cleanup efforts. EPA START will be performing an SPCC inspection of the facility on 12/13/2011.

During the period of 12/12/2011 through 12/14/2011, LP continued flushing operations and have flushed approximately 0.30 miles of the 0.5 miles of impacted creek area. During flushing activities, approximately 1,400 BBL of oil/salt water/water was recovered. Out of the 1,400 BBL recovered, an estimated 4 BBL is oil. LP continued the removal of oiled sediment and vegetation along the creek. LP continued excavating the soil around the creek entrance utilizing shovels due to the multiple flow lines in the area. LP utilized absorbent pads to collect residual oil after the flushing operations. LP retained an additional crew for cleanup activities. LP will reduce the operational period from 24 hours a day to roughly 16 hours a day. EPA START was on site documenting cleanup activities. EPA START conducted an SPCC inspection of the tank battery. OSC Brescia issued a formal Notice of Federal Interest(NOFI) letter to LP. EPA START demobilized on the evening of 12/14/2011. LP will provide the OSC daily cleanup reports until the SOW has been completed. .

During the period of 12/14/2011 through 12/17/2011, LP continued flushing operations and completed the remaining 0.20 miles of impacted creek. During flushing activties, approximately 400 BBL of oil/salt water/water was recovered. Out of the 400 BBL recovered an estimated 0.5 BBL is oil. LP utilized absorbent pads in the creek area to remove any residual oil after the flushing operations. LP completed excavating impacted soil around the creek entrance. LP placed additional absorbent boom into the creek at several locations to collect any residual oil that is discharged from future rain events. LP contacted OSC Brescia and requested a site walk of the creek. LP is currently conducting pad/boom maintenance in the creek as needed. OSC Brescia is scheduled to conduct a site walk during the week of Dec 19-23, 2011.

On 12/20/2011 OSC Brescia and OSC Todd arrived on site along with TXPW and conducted a site walk with LP. The site walk included walking the 0.5 miles of impacted creek and the area where the oil impacted soil had been excavated. Upon inspection, LP had placed multiple sets of absorbent boom within the creek area to recover any residual oil if a rain event occurred. LP informed OSC Brescia that they would continue to provide maintenance as needed on the boom and by utilizing pads to collect any residual oil that became present over the next few weeks. The oil impacted soil had been stockpiled on site and LP is currently following TRRC regulations on future soil use. Clean soil will be placed in the excavated areas and will be seeded. All waste generated utilizing pads and boom will be disposed of properly. All oil/salt water/rinsate water collected during cleanup operations was disposed of on site through injection wells.  OSC Brescia determined that the agreed upon SOW had been completed and that the impacted creek had been properly cleaned. LP informed OSC Brescia that they would submit a final report to close out the emergency removal action. OSC Brescia, OSC Todd and TXPW demobilized from the site.


2.1.3 Enforcement Activities, Identity of Potentially Responsible Parties (PRPs)

2.1.4 Progress Metrics

Waste Stream Medium Quantity Manifest # Treatment Disposal
           
           
           


  2.2 Planning Section
    2.2.1 Anticipated Activities
LP will continue maintenance of booms/pads in creek as needed to recover any residual oil from future rain events.

2.2.1.1 Planned Response Activities
No further action is needed.

2.2.1.2 Next Steps
No further action is needed.

2.2.2 Issues
None

  2.3 Logistics Section
    No information available at this time.

  2.4 Finance Section
    No information available at this time.

  2.5 Other Command Staff
    No information available at this time.

3. Participating Entities
  No information available at this time.

4. Personnel On Site
  No information available at this time.

5. Definition of Terms
  No information available at this time.

6. Additional sources of information
  No information available at this time.

7. Situational Reference Materials
  No information available at this time.